# ForeSight: A Predictive-Scheduling Deterministic Database

### v1

@@ 7.4. Effectiveness of Deterministic Reordering
In this section, we evaluate the effectiveness of the deterministic reordering algorithm MTFS introduced in Section 6. The evaluation compares two implementations of MTFS —single-threaded baseline (ForeSight-S) and the fully parallel implementation (ForeSight-M)—against representative systems with reorder algorithm, including Aria (Lu et al., 2020) and FGA (Ding et al., 2018). This evaluation allows us to highlight the benefits of predictive scheduling under different execution environments and workload characteristics.
@@ 8. Related Work
Deterministic Database. The concept of deterministic database systems can be traced back to the late 1990s (Thomson and Abadi, 2010b; Narasimhai et al., 1999; Jiménez-Peris et al., 2000; Abadi and Faleiro, 2018). The central idea is to enforce consistent execution across replicas by producing identical thread interleavings through a deterministic scheduler (Ren et al., 2012, 2014, 2015; Faleiro et al., 2014; Thomson and Abadi, 2010a). Representative systems such as Calvin (Thomson et al., 2012), BOHM (Faleiro and Abadi, 2015), and PWV (Faleiro et al., 2017) adopt this approach by constructing dependency graphs or ordered locks from pre-obtained read/write sets. These techniques guarantee deterministic execution but restrict flexibility for workloads with data-dependent or interactive transactions. To avoid dependency tracking and locking, Aria (Lu et al., 2020) instead executes batches in parallel on consistent snapshots and deterministically validates them at commit time. Later work includes AriaER (Miyazaki et al., 2024), which reduces aborts via early elimination of write–write conflicts, and DMUCCA (Li et al., 2024), which reorders batches using access-pattern weights to improve throughput. Despite these advances, existing deterministic databases still face challenges such as the overhead of dependency tracking and delayed conflict resolution, motivating the need for approaches that remain efficient under high contention.
@@ 9. Conclusion
In this paper, we introduced ForeSight, a deterministic database system that achieves high performance through predictive, conflict-aware scheduling. By integrating lightweight conflict prediction, multi-version-based optimization, and efficient reordering, ForeSight reduces aborts and improves throughput without requiring prior knowledge of transaction read/write sets. Extensive Evaluations on standard benchmarks confirm its effectiveness in handling high-contention workloads while maintaining deterministic guarantees. By introducing foresight into scheduling, ForeSight charts a promising direction for scalable, reliable, and efficient deterministic databases. Additionally, future work will explore further optimizations in conflict prediction and scheduling strategies to enhance performance even more.

### v2

@@ 7.4. Effectiveness of Deterministic Reordering
In this section, we evaluate the effectiveness of the deterministic reordering algorithm MTFS introduced in Section 6. The evaluation compares two implementations of MTFS, the single-threaded baseline ForeSight-S and the fully parallel implementation ForeSight-M, against representative systems with reorder algorithms, including Aria (Lu et al., 2020) and FGA (Ding et al., 2018). This evaluation allows us to highlight the benefits of predictive scheduling under different execution environments and workload characteristics.
@@ 8. Related Work
Deterministic Database. The concept of deterministic database systems can be traced back to the late 1990s (Thomson and Abadi, 2010b; Narasimhai et al., 1999; Jiménez-Peris et al., 2000; Abadi and Faleiro, 2018). The central idea is to enforce consistent execution across replicas by producing identical thread schedules through a deterministic scheduler (Ren et al., 2012, 2014, 2015; Faleiro et al., 2014; Thomson and Abadi, 2010a). Representative systems such as Calvin (Thomson et al., 2012), BOHM (Faleiro and Abadi, 2015), and PWV (Faleiro et al., 2017) adopt this approach by constructing dependency graphs or ordered locks from pre-obtained read/write sets. These techniques guarantee deterministic execution but restrict flexibility for workloads with data-dependent or interactive transactions. To avoid the overhead of dependency analysis and locking, Aria (Lu et al., 2020) instead executes batches in parallel on consistent snapshots and deterministically validates them at commit time. Later work includes AriaER (Miyazaki et al., 2024), which reduces aborts via early elimination of write-write conflicts, and DMUCCA (Li et al., 2024), which reorders batches using access-pattern weights to improve throughput. However, existing deterministic databases still face challenges such as the overhead of dependency analysis and conflict resolution, motivating the need for approaches that remain efficient under high contention.
@@ 9. Conclusion
In this paper, we introduced ForeSight, a deterministic database system that achieves high performance through predictive, conflict-aware scheduling. By integrating lightweight conflict prediction, multi-version-based optimization, and efficient reordering, ForeSight reduces aborts and improves throughput without requiring prior knowledge of transaction read/write sets. Extensive Evaluations on standard benchmarks confirm its effectiveness in handling high-contention workloads while maintaining deterministic guarantees. By introducing foresight into scheduling, ForeSight offers a promising direction for scalable, reliable, and efficient deterministic databases. Additionally, future work will explore further optimizations in conflict prediction and scheduling strategies to enhance performance even more.